# Ringdown amplitudes of nonspinning eccentric binaries

### v1

@@ Abstract
Closed-form expressions for the ringdown complex amplitudes of nonspinning unequal-mass binaries in arbitrarily eccentric orbits are presented. They are built upon 237 numerical simulations contained within the RIT catalog, through the parameterisation introduced in [Phys. Rev. Lett. 132, 101401]. Global fits for the complex amplitudes, associated to linear quasinormal mode frequencies of the dominant ringdown modes, are obtained in a factorised form immediately applicable to any existing quasi-circular model. Similarly to merger amplitudes, ringdown ones increase by more than 50% compared to the circular case for high impact parameters (medium eccentricities), while strongly suppressed in the low impact parameter (highly eccentric) limit. Such reduction can be explained by a transition between an “orbital-type” and an “infall-type” dynamics. The amplitudes (phases) fits accuracy lies around a few percent (deciradians) for the majority of the dataset, comparable to the accuracy of current state-of-the-art quasi-circular ringdown models, and well within current statistical errors of current LIGO-Virgo-Kagra ringdown observations. These expressions constitute another building block towards the construction of complete general-relativistic inspiral-merger-ringdown semi-analytical templates, and allow to extend numerically-informed spectroscopic analyses beyond the circular limit. Such generalisations are key to achieve accurate inference of compact binaries astrophysical properties, and tame astrophysical systematics within observational investigations of strong-field general relativistic dynamics.
@@ III Waveform model
Finally, $t_{\rm ref}$ represents the time at which complex amplitude parameters are defined. Given the dependence of $\omega_{\ell mn}$ on the BH spin, different choices of $t_{\rm ref}$ rescale the amplitude factors of different binary configurations in a non-trivial way. In Ref. [34], it was found that setting reference time $t_{\rm ref}$ to the maximum of $|h_{22}|$, $t_{\rm ref}=t_{\rm peak}$, allows to simplify the complex amplitude dependence on binary parameters. Here, the same logic is followed, while accounting for the fact that in highly eccentric scenarios the peak of the waveform does not in general provide a good estimate of the merger time. This is because for intermediate eccentricity values, the last periastron before merger can display a larger amplitude compared to the merger itself, as displayed e.g. in Fig. 5. See also Figs. 1, 6 of Ref. [48] or Supplementary Fig. 3 of Ref. [24], diplaying conceptually identical phenomenologies. Hence, the definition $t_{\rm ref}=t_{\rm mrg}$ is used, where $t_{\rm mrg}$ is the last peak of $A_{22}$ immediately before a ringdown regime begins, an appropriate estimate of the merger time. For an extended discussion of this point, see Ref. [18]. This choice allows to automatically obtain peak-rescaled amplitude values when varying the fit start time. The above $t_{\rm ref}$ parameter should not be confused with $t_{\rm start}$ [^3] Differently from QNM, which are excited during light-ring crossing, tails are mostly excited during apastra [46], hence will have in general a different starting time. Since tail terms will not be considered here, this point will not enter the fitting procedure below, hence for ease of notation $t_{\rm start}=t^{\rm ring}_{\rm start}$, setting the validity regime of a model built through a superposition of QNMs with constant mass, spin and complex amplitude. Earlier times are contaminated by transients [49, 50, 45], non-linearities [51, 52, 53, 54, 55, 50, 44, 54, 56, 57, 58, 59, 60, 61] and mass-spin variations [53, 62, 63, 64], for which no first-principles model exists at present, with current templates relying on phenomenological constructions [65, 66, 67, 68, 10, 69, 70, 71, 72, 73]. Hence, the scope of the present investigation will be limited to construct a model for the late-time constant amplitudes, possessing a robust theoretical base and a clean physical interpretation.
@@ IV.2 Extraction procedure
The fit is performed using a Bayesian method. The algorithm, implemented in the bayRing package [78], is described in detail in Ref. [58] and relies on the cpnest sampling algorithm [79]. Given the simple inference problem at hand, 32 live points and maximum Markov Chain Monte Carlo steps, with a single chain, are used. Such sampler settings were tested to be sufficient to obtain an unbiased extraction, as the results did not change when increasing sampler settings. Wide parameters ranges are explored, with priors uniform in: $[0,2\pi]$ on phase parameters, $[-20,5]$ on $ln\,A^{\pm}$ and $ln\,A^{T}$, for all the modes.
@@
The model includes only the dominant $(\ell,m,0)$ fundamental mode for each of the the $(\ell,m)=[(2,2),(3,3),(2,1),(4,4)]$ spherical harmonics considered. Attempts to construct a robust global fit the overtones, mode-mixing and counter-rotating contributions, expected to constitute the next dominant component, failed due to the limited available data quality. Similarly, a robust fit of the $(\ell,m,n)=(2,0,0)$ mode could not be obtained. This is expected since we do not include eccentricity values close to unity, which is the regime where this mode is expected to be strongly excited, and since memory contributions (dominant for such mode [80]) are not accurately resolved in the dataset under consideration.
@@ V.2 Fit construction
First, instead of modeling the amplitudes (phases) directly, their ratio (difference) is modeled compared to the quasi-circular case. This choice allows to scale any leading order dependence (e.g. on $\nu$ or $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$) observed independently of the binary orbital class. It also allows to obtain a factorized (additive) form for our fits, which can be straightforwardly applied on top of quasi-circular values [91, 92, 52, 34, 87, 47], maintaining the accuracy of the quasi-circular limit where more numerical simulations of higher accuracy are available. Specifically, for $(\ell,m,n)=[(2,2,0),(3,3,0),(2,1,0),(4,4,0)]$, the following quantities are modeled:
@@
Quasi-circular values obtained in Ref. [47] are used, since this most recent fit relies on the highly-accurate SXS catalogue [93]. The additional subtraction of the $m\cdot\phi_{220}$ factor serves to eliminate the dependence on the arbitrary phase of the specific simulation, similarly to Refs. [39, 87, 47]. This is not a restrictive assumption, since it amounts to a global phase that can always be reabsorbed in the spherical-harmonic $e^{im\phi}$ factor, which explains the $m$ prefactor applied in the subtraction. The above definition automatically imposes $\phi_{220}=0$, consistently with the conventions of Ref. [47], which is followed here. The phases $\hat{\phi}_{\ell mn}$ are also further unwrapped, i.e. the transformation $\hat{\phi}_{\ell mn}\rightarrow\hat{\phi}_{\ell mn}+2n\pi$ is applied for the smallest possible integer $n$ that ensures continuity. Above, a previous unwrapping stage was applied, ensuring continuity only at the level of a single binary extraction. This second unwrapping stage is instead required to ensure continuity at the level of the entire dataset. Since now the phase values to unwrap are multi-dimensional (as opposed to the one-dimensional unwrapping stage applied above), the numpy.unwrap function [94] serves to enforce continuity for most regions of the parameter space, but not for the entire set of datapoints of all modes, since the numpy unwrapping algorithm is constructed for one-dimensional functions. More sophisticated multi-dimensional unwrapping procedures should in principle have been applied, but this step was found to be unnecessary, as continuity could be enforced by manual addition of $2\pi$ factors for the $(2,1,0),(4,4,0)$ modes in regions of the parameter space easy identified in three-dimensional plots  [^4] A multi-dimensional unwrapping algorithm is available e.g. at github.com/geggo/phase-unwrap.. The resulting phases are then continuous, although now not in general within the $[0,2\pi]$ interval, which is still allowed given the phases periodicity.
@@
where $(Y_{0},\,p_{k,i})\in\mathbb{R}$, $N$ is the number of fitting variables considered. In practise, $N=2$ for “quasi-universal” fits for which $Q_{i}$ runs on $(\nu,\,\hat{b}_{\mathrm{mrg}})$, while $N=3$ in the cases where $Q_{i}$ runs on $(\nu,\,\hat{b}_{\mathrm{mrg}},\,\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}})$. The $Y$ variable denotes either of the fit quantities: $(\hat{A}_{\ell mn},\,\hat{\phi}_{\ell mn})$ for $(\ell,m,n)=[(2,2,0),(3,3,0),(2,1,0),(4,4,0)]$. This class of functions was found to be sufficiently simple and flexible to capture the structure of the complex amplitudes considered here, similarly to what observed in Ref. [18].
@@
Residuals for the amplitudes are defined by $\Delta Y\equiv(Y-Y_{\mathrm{NR}})/Y_{\mathrm{NR}}$, while for the phases it is more meaninful to report residuals as $\Delta Y\equiv(Y-Y_{\mathrm{NR}})$, to avoid discontinuities due to zero-crossings. The obtained coefficients and a python implementation of the fits will be made publicly available at: github.com/GCArullo/noncircular_BBH_fits.
@@ V.3 Results
Results for the equal-mass case are displayed in Fig. 4 for the dominant quadrupolar mode. The top panel reports on the one-dimensional fit in terms of a single noncircular variable $\hat{b}_{\mathrm{mrg}}$, which was found to be highly effective in modeling merger-remnant quantities in Ref. [18]. There, quasi-universality held at the level of a few percent for most of the cases considered in Ref. [18]. Here, a spiral-like structure in the residuals indicates the degree of such universality violation, and points to the need of adding a second variable to the fit. The two-dimensional case, now in terms of both $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$ and $\hat{b}_{\mathrm{mrg}}$ is displayed in the bottom panel. The two-dimensional fit diplays a smooth hyperboloidal structure, and residuals decrease below $10\%$, with the exception of a few outliers. Specifically, $92\%$ of the simulations in the dataset have residuals smaller than $5\%$, and $97\%$ below $10\%$. Similar considerations apply to the dominant hexadecapolar mode (odd-$m$ modes are zero by symmetry in the equal-mass case [33]), as displayed in Fig. 8. The amplitude fit accuracy remains very similar to the one of the quadrupolar mode. In this case we can also fit for the phase parameter (in the quadrupolar case, the phase is zero by construction), and similar fits give residuals below $\rm 0.2\,rad$ (except for a single case). This is comparable to the accuracy of existing quasi-circular models, typically around a few percent in amplitude and a few deciradians in phase [87].
@@
The scatter and large residuals observed in the $\hat{b}_{\mathrm{mrg}}$ fits point to the requirement of employing the full dimensionality of the noncircular initial conditions. Hence, Fig. 7 presents the residuals of the fit constructed in terms of $(\nu,\hat{b}_{\mathrm{mrg}},\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}})$. Agreement improves considerably, with $87\%$ of the simulations displaying an error smaller than $5\%$ on the noncircular correction to the $A_{220}$ amplitude, with $97.5\%$ smaller than $10\%$. These numbers decrease to $85\%$ ($82\%$) below $5\%$ and $95\%$ ($93\%$) below $10\%$ for the correction to $A_{330}$ ($A_{210}$). The $\phi_{330}$ parameter displays high accuracy, with $97\%$ of the residuals below $\rm 0.2\,rad$, and only one simulation above $\rm 0.5\,rad$. The accuracy slightly decreases for $\phi_{210}$, for which only $85\%$ ($96\%$) of the simulations have residuals below $\rm 0.2\,rad$ ($\rm 0.5\,rad$). Apart from a few percent outliers, such level of accuracy is comparable to the one of state-of-the-art quasi-circular models, typically around the percent-level in amplitudes and $\rm 0.2/0.3\,rad$ in phase, see Ref. [87]. Fit results and residuals for the $(4,4,0)$ mode amplitudes and phases are shown in Fig. 9, and are similar to the ones described above, albeit less accurate as expected.
These results robustly indicate that ringdown amplitudes can increase by more than 50% compared to the quasi-circular case for high values of the impact parameter (medium eccentricities). A few datapoints indicate that such increase might reach even 100%, albeit these appear as outliers compared to the bulk of the dataset. Additional high-quality simulations are needed to assess the robustness of such even larger increase. The low impact parameter (highly eccentric) limit is instead consistently represented by the fit, with a suppression of up to 50%.

### v2

@@ Abstract
Closed-form expressions for the ringdown complex amplitudes of nonspinning unequal-mass binaries in arbitrarily eccentric orbits are presented. They are built upon 237 numerical simulations contained within the RIT catalog, through the parameterisation introduced in [Phys. Rev. Lett. 132, 101401]. Global fits for the complex amplitudes, associated to linear quasinormal mode frequencies of the dominant ringdown modes, are obtained in a factorised form immediately applicable to any existing quasi-circular model. Similarly to merger amplitudes, ringdown ones increase by more than 50% compared to the circular case for high impact parameters (medium eccentricities), while strongly suppressed in the low impact parameter (highly eccentric) limit. Such reduction can be explained by a transition between an “orbital-type” and an “infall-type” dynamics. The amplitudes (phases) fits accuracy lies around a few percent (deciradians) for the majority of the dataset, comparable to the accuracy of current state-of-the-art quasi-circular ringdown models, and well within current statistical errors of current LIGO-Virgo-Kagra ringdown observations. These expressions constitute another building block towards the construction of complete general relativistic inspiral-merger-ringdown semi-analytical templates, and allow to extend numerically-informed spectroscopic analyses beyond the circular limit. Such generalisations are key to achieve accurate inference of compact binaries astrophysical properties, and tame astrophysical systematics within observational investigations of strong-field general relativistic dynamics.
@@ III Waveform model
Finally, $t_{\rm ref}$ represents the time at which complex amplitude parameters are defined. Given the dependence of $\omega_{\ell mn}$ on the BH spin, different choices of $t_{\rm ref}$ rescale the amplitude factors of different binary configurations in a non-trivial way. In Ref. [34], it was found that setting reference time $t_{\rm ref}$ to the maximum of $|h_{22}|$, $t_{\rm ref}=t_{\rm peak}$, allows to simplify the complex amplitude dependence on binary parameters. Here, the same logic is followed, while accounting for the fact that in highly eccentric scenarios the peak of the waveform does not in general provide a good estimate of the merger time. This is because for intermediate eccentricity values, the last periastron before merger can display a larger amplitude compared to the merger itself, as displayed e.g. in Fig. 5. See also Figs. 1, 6 of Ref. [49] or Supplementary Fig. 3 of Ref. [24], diplaying conceptually identical phenomenologies. Hence, the definition $t_{\rm ref}=t_{\rm mrg}$ is used, where $t_{\rm mrg}$ is the last peak of $A_{22}$ immediately before a ringdown regime begins, an appropriate estimate of the merger time. For an extended discussion of this point, see Ref. [18]. This choice allows to automatically obtain peak-rescaled amplitude values when varying the fit start time. The above $t_{\rm ref}$ parameter should not be confused with $t_{\rm start}$ [^3] Differently from QNM, which are excited during light-ring crossing, tails are mostly excited during apastra [46], hence will have in general a different starting time. Since tail terms will not be considered here, this point will not enter the fitting procedure below, hence for ease of notation $t_{\rm start}=t^{\rm ring}_{\rm start}$, setting the validity regime of a model built through a superposition of QNMs with constant mass, spin and complex amplitude. Earlier times are contaminated by transients [50, 51, 45], non-linearities [52, 53, 54, 55, 56, 51, 44, 55, 57, 58, 59, 60, 61, 62] and mass-spin variations [54, 63, 64, 65, 66], for which no first-principles model exists at present, with current templates relying on phenomenological constructions [67, 68, 69, 70, 10, 71, 72, 73, 74, 75]. Hence, the scope of the present investigation will be limited to construct a model for the late-time constant amplitudes, possessing a robust theoretical base and a clean physical interpretation.
@@ IV.2 Extraction procedure
The fit is performed using a Bayesian method. The algorithm, implemented in the bayRing package [80], is described in detail in Ref. [59] and relies on the cpnest sampling algorithm [81]. Given the simple inference problem at hand, 32 live points and maximum Markov Chain Monte Carlo steps, with a single chain, are used. This values are increased to 64 when considering mode-mixing contributions, which double the number of free parameters. Such sampler settings were tested to be sufficient to obtain an unbiased extraction, as the results did not change when increasing sampler settings. Wide parameters ranges are explored, with priors uniform in: $[0,2\pi]$ on phase parameters, $[-20,5]$ on $ln\,A^{\pm}$ and $ln\,A^{T}$, for all the modes.
@@
The model includes only the dominant $(\ell,m,0)$ fundamental mode for each of the $(\ell,m)=[(2,2),(3,3),(2,1),(4,4)]$ spherical harmonics considered. For the $(\ell,m)=(3,2)$ harmonic, the dominant mode-mixing contribution from the $(\ell,m,n)=(2,2,0)$ mode was also included, in addition to the $(3,2,0)$. Attempts to construct a robust global fit the overtones and counter-rotating contributions, expected to constitute the next dominant component, failed due to the limited available data quality. Similarly, a robust fit of the $(\ell,m,n)=(2,0,0)$ mode could not be obtained. This is expected since we do not include eccentricity values close to unity, which is the regime where this mode is expected to be strongly excited, and since memory contributions (dominant for such mode [82]) are not accurately resolved in the dataset under consideration.
@@ V.2 Fit construction
First, instead of modeling the amplitudes (phases) directly, their ratio (difference) is modeled compared to the quasi-circular case. This choice allows to scale any leading order dependence (e.g. on $\nu$ or $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$) observed independently of the binary orbital class. It also allows to obtain a factorized (additive) form for our fits, which can be straightforwardly applied on top of quasi-circular values [93, 94, 53, 34, 89, 48], maintaining the accuracy of the quasi-circular limit where more numerical simulations of higher accuracy are available. Specifically, for $(\ell,m,n)=[(2,2,0),(3,3,0),(2,1,0),(4,4,0),(3,2,0)]$, the following quantities are modeled:
@@
Quasi-circular values obtained in Ref. [48] are used, since this most recent fit relies on the highly-accurate SXS catalogue [95]. The additional subtraction of the $m\cdot\phi_{220}$ factor serves to eliminate the dependence on the arbitrary phase of the specific simulation, similarly to Refs. [39, 89, 48]. This is not a restrictive assumption, since it amounts to a global phase that can always be reabsorbed in the spherical-harmonic $e^{im\phi}$ factor, which explains the $m$ prefactor applied in the subtraction. The above definition automatically imposes $\phi_{220}=0$, consistently with the conventions of Ref. [48], which is followed here. The phases $\hat{\phi}_{\ell mn}$ are also further unwrapped, i.e. the transformation $\hat{\phi}_{\ell mn}\rightarrow\hat{\phi}_{\ell mn}+2n\pi$ is applied for the smallest possible integer $n$ that ensures continuity. Above, a previous unwrapping stage was applied, ensuring continuity only at the level of a single binary extraction. This second unwrapping stage is instead required to ensure continuity at the level of the entire dataset. Since now the phase values to unwrap are multi-dimensional (as opposed to the one-dimensional unwrapping stage applied above), the standard numpy.unwrap function [96] serves to enforce continuity for most regions of the parameter space, but not for the entire set of datapoints of all modes. This is because the numpy unwrapping algorithm is constructed for one-dimensional functions. More sophisticated multi-dimensional unwrapping procedures should in principle have been applied, but this step was found to be unnecessary, as continuity could be enforced by manual addition of $2\pi$ factors for the $(2,1,0),(4,4,0),(3,2,0)$ modes in regions of the parameter space easy identified in three-dimensional plots  [^4] A multi-dimensional unwrapping algorithm is available e.g. at github.com/geggo/phase-unwrap.. The resulting phases are then continuous, although now not in general within the $[0,2\pi]$ interval, which is still allowed given the phases periodicity.
@@
where $(Y_{0},\,p_{k,i})\in\mathbb{R}$, $N$ is the number of fitting variables considered. In practise, $N=2$ for “quasi-universal” fits for which $Q_{i}$ runs on $(\nu,\,\hat{b}_{\mathrm{mrg}})$, while $N=3$ in the cases where $Q_{i}$ runs on $(\nu,\,\hat{b}_{\mathrm{mrg}},\,\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}})$. The $Y$ variable denotes either of the fit quantities: $(\hat{A}_{\ell mn},\,\hat{\phi}_{\ell mn})$ for $(\ell,m,n)=[(2,2,0),(3,3,0),(2,1,0),(4,4,0),(3,2,0)]$. This class of functions was found to be sufficiently simple and flexible to capture the structure of the complex amplitudes considered here, similarly to what observed in Ref. [18].
@@
Residuals for the amplitudes are defined by $\Delta Y\equiv(Y-Y_{\mathrm{NR}})/Y_{\mathrm{NR}}$, while for the phases it is more meaninful to report residuals as $\Delta Y\equiv(Y-Y_{\mathrm{NR}})$, to avoid discontinuities due to zero-crossings. The obtained coefficients and a python implementation of the fits are publicly available at: github.com/GCArullo/noncircular_BBH_fits.
@@ V.3 Results
Results for the equal-mass case are displayed in Fig. 4 for the dominant quadrupolar mode. The top panel reports on the one-dimensional fit in terms of a single noncircular variable $\hat{b}_{\mathrm{mrg}}$, which was found to be highly effective in modeling merger-remnant quantities in Ref. [18]. There, quasi-universality held at the level of a few percent for most of the cases considered in Ref. [18]. Here, a spiral-like structure in the residuals indicates the degree of such universality violation, and points to the need of adding a second variable to the fit. The two-dimensional case, now in terms of both $\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}}$ and $\hat{b}_{\mathrm{mrg}}$ is displayed in the bottom panel. The two-dimensional fit diplays a smooth hyperboloidal structure, and residuals decrease below $10\%$, with the exception of a few outliers. Specifically, $92\%$ of the simulations in the dataset have residuals smaller than $5\%$, and $97\%$ below $10\%$. Similar considerations apply to the dominant hexadecapolar mode and to the dominant mode receiving mode-mixing, $(\ell,m)=(3,2)$, as displayed in Figs. 8,9. Odd-$m$ modes are zero by symmetry in the equal-mass case [33]. The amplitude fit accuracy remains similar to the one of the quadrupolar mode. For these modes we can also fit for the phase parameter (in the (2,2,0) case, the phase is zero by construction), and similar fits give residuals below $\rm 0.2\,rad$ (except for a single case) for the $(4,4,0)$ mode, while the $(3,2,0)$ case displays a few outliers above $\rm 0.5\,rad$. Visual inspection of the underlying waveforms suggest the (3,2,0) outliers are driven by systematic errors in the numerical data, and point to the need of improved numerical simulations. Except for the few cases just discussed, the equal mass fits are however comparable to the accuracy of existing quasi-circular models, typically around a few percent in amplitude and a few deciradians in phase [89].
@@
The scatter and large residuals observed in the $\hat{b}_{\mathrm{mrg}}$ fits point to the requirement of employing the full dimensionality of the noncircular initial conditions. Hence, Fig. 7 presents the residuals of the fit constructed in terms of $(\nu,\hat{b}_{\mathrm{mrg}},\hat{E}_{\mathrm{eff}}^{\mathrm{mrg}})$. Agreement improves considerably, with $87\%$ of the simulations displaying an error smaller than $5\%$ on the noncircular correction to the $A_{220}$ amplitude, with $97.5\%$ smaller than $10\%$. These numbers decrease to $85\%$ ($82\%$) below $5\%$ and $95\%$ ($93\%$) below $10\%$ for the correction to $A_{330}$ ($A_{210}$). The $\phi_{330}$ parameter displays high accuracy, with $97\%$ of the residuals below $\rm 0.2\,rad$, and only one simulation above $\rm 0.5\,rad$. The accuracy slightly decreases for $\phi_{210}$, for which only $85\%$ ($96\%$) of the simulations have residuals below $\rm 0.2\,rad$ ($\rm 0.5\,rad$). Apart from a few percent outliers, such level of accuracy is comparable to the one of state-of-the-art quasi-circular models, typically around the percent-level in amplitudes and $\rm 0.2/0.3\,rad$ in phase, see Ref. [89]. Fit results and residuals for the $(4,4,0)$ mode amplitudes and phases are shown in Fig. 10, and are similar to the ones described above, albeit less accurate as expected. Instead, Fig. 11 displays the results for the dominant harmonic receiving mode-mixing, $(\ell,m)=(3,2)$. As discussed in the equal-mass results description, systematic errors in the numerical data seem to affect this mode to a larger degree, yielding a few outliers above the 20% ($\rm 0.5\,rad$) level for the amplitude (phase) parameters. However, residuals for the majority of the simulations still present errors of the same order of magnitude as other modes, with $91\%$ of the dataset carrying residuals below $10\%$ ($\rm 0.2rad$) for the amplitude (phase) parameter. The mass ratio dependence presents a richer structure in this case, with a non-monotonic behaviour in the amplitude and a large drop in the phase around $\nu=0.21$. Both these features also appear in the quasi-circular case [53].
In summary, fit results robustly indicate that ringdown amplitudes can increase by more than 50% compared to the quasi-circular case for high values of the impact parameter (medium eccentricities). A few datapoints indicate that such increase might reach even 100%, albeit these appear as outliers compared to the bulk of the dataset. Additional high-quality simulations are needed to assess the robustness of such even larger increase. The low impact parameter (highly eccentric) limit is instead smoothly represented by the fit, with a suppression of up to 50%.